22:30 Shift handed over. DAQ OK. Target OK. Beam OK.
00:22 Temperatures OK
Stats OK
Good events OK
Scalers OK
System wide checks as usual - WR fails by 68 ns
Leakage current OK
Target OK
Motors OK
Pressure in ring 5E-10 mbar OK
Beam OK
02:26 - All OK as above
WR now fails by 80
Data rate ~800 KB/s
03:40 All OK as above
Data rate back to ~ 1900 KB/s
04:20 All OK as above
Started compressing files in WS3, monitor 1
Noticed bias is not as stable as it has been so far, but no reason for concern. See attach.
04:30 Saw a spike in counts similar to beam halo. See attachment. Did not see it before. No effects on leakage current. To be checked offline.
06:00 Could not see any more spikes
All other checks OK as before.
Data rate back to ~1000 Kb/s
06:32 Requested beam off for end of beamtime.
Left DAQ on - run R88
Turning off jet target as per Nikos instructions
Final cleanup + pulse walkthrough to be done by JM and OH later today.
Target nozzle removed - pressure in S1 now 1.93E-8 mbar
Valves closed
Entering cave to switch compressor off and change N backing pressure to 10 bar. |